56/6+2/1=7 Insert paranthesis to make the answer equal zero
Answer by josmiceli(9661)

56/6+2/1=7
56/(6+2)/1 = 7
56/8/1 = 7
division by 1 doesn't change anything, so cross out the 1
7 = 7